The company purchased and restored a classic 1957 Thunderbird convertible 7 years ago for $8,300.00. Today at auction, the car sold for $83,000.00.


Holding period and annual (Investment) returns. Bohenick Classic Automobiles restores and rebuilds old classic cars. The company purchased and restored a classic 1957 Thunderbird convertible 7 years ago for $8,300.00. Today at auction, the car sold for $83,000.00. What are the holding period return and the annual return on this investment? 


What is the holding period return of the car?
Answer #1

Holding period return = ((selling price / purchase price) - 1) * 100
= (($83,000 / $8,300) - 1) * 100
= 900%

Annual return on this investment = (1 + HPR)^(1/n) - 1
= (1 + 900%)^^(1/7) - 1
= 1.3895 - 1
= 38.95%